# 3D Foundation AI Model for Generalizable Disease Detection in Head Computed Tomography

### v1

@@ Appendix B Data augmentation details
We applied Random Flipping across all three dimensions, Random Shift Intensity with offset $0.1$ for both pre-training and fine-tuning. For DINO training. random Gaussian Smoothing with sigma=$(0.5-1.0)$ is applied across all dimensions, Random Gamma Adjust is applied with gamma=$(0.2-1.0)$.
Appendix C Additional experiment results
This section provides additional experimental results with more details. Supplementary Figs. 3 and 4 compares the performance of the foundation model using different numbers of channels and patch sizes, demonstrating that the architecture design of our foundation model is optimal.
Supplementary Fig. 5 compares our foundation model with a foundation CT model from previous studies, Merlin[20], which was trained on abdomen CT scans with corresponding radiology report pairs. Our model demonstrates superior performance on head CT scans.
Supplementary Fig. 7 compares our foundation model with Google CT Foundation model [17], which was trained on large scale and diverse CT scans from different anatomy with corresponding radiology report pairs. Our model consistently shows improved performance across the board even though our model was pre-trained with less samples.
Supplementary Fig. 6 compares the performance on downstream tasks with various supervised tuning methods applied to foundation models pretrained with the MAE and DINO frameworks. Per-pathology comparisons are shown in Supplementary Figs. 9 and 8. Meanwhile, supplementary Fig. 10 complements Fig. 3, illustrating the per-pathology performances of foundation models pretrained with different scales of training data.
Supplementary Figs. 11 and 12 studies the impact of batch effect caused by different CT scan protocols of slice thickness and machine manufacturer. Detailed per-pathology performances are shown in Supplementary Figs. 13 and 14.
@@ Appendix C Additional experiment results
Supplementary Figure 5: Comparison to previous 3D Foundation Model. This plot compares the performance of our model with Merlin [20] and models trained from scratch across four datasets for our model and ResNet50-3D. Our DINO trained model is used in this comparison. Our model demonstrates consistently superior performance across majority of diseases, with the exception of epidural hemorrhage (EDH) in the CQ500 dataset.
@@
Supplementary Figure 7: Performance comparison of linear probing for Our Model vs. Google CT Foundation model This plot compares our model performance vs. Google CT Foundation model[38] and Merlin [20] across all diseases on RSNA and CQ500. Since Google CT Foundation moudel requires uploading data to Google Cloud (not allowed on our private data) for requesting model embeddings with model weights inaccessible, only public dataset comparison is provided in this study. Similar to other evaluations, we observed that our model outperforms Google CT Foundation model across the board with the only exception on Midline Shift for Google CT Foundation model and EDH for Merlin.
@@
Supplementary Figure 8: Performance comparison of supervised finetuning methods per pathology on the foundation model trained with DINO. This plot breaks down the average performance across all diseases shown in Supplementary Fig. 6. The results show that fine-tuning the entire network achieves the best performance in most scenarios. However, linear probing closely approaches finetuning performance for many diseases especially on small or imbalanced dataset, underscoring the capability of our pre-trained models to generate representations that adapt effectively to diverse disease detection tasks.
@@
Supplementary Figure 10: Performance for Different Percentage of Pre-training Samples (Per-Pathology). This plot illustrates label efficiency for individual pathologies using Tukey plots, alongside the average performance across all diseases shown in Fig. 3. The results indicate that the majority of pathologies show improved downstream performance as the amount of pretraining data increases.
Appendix D Time complexity increase with reduced patch size
@@
Supplementary Figure 11: The 2D projection with t-SNE of CT volume representation extracted from the foundation model. Interestingly, certain subgroups still exhibited slightly better AUCs. For instance, scans with slice thicknesses between 1–4 mm (represented by light green points in the upper panel of Fig. 11) align with a specialized protocol for CT angiography (CTA), which uses contrast dye to improve diagnosis on particular diseases.

### v2

@@ Appendix B Data augmentation details
We applied Random Flipping across all three dimensions, Random Shift Intensity with offset $0.1$ for both pre-training and fine-tuning. For DINO training. random Gaussian Smoothing with sigma=(0.51.0)(0.5-1.0) is applied across all dimensions, Random Gamma Adjust is applied with gamma=(0.21.0)(0.2-1.0).
Appendix C Clarification on model evaluation
While previous literature achieved promising performance on Head CT disease detection tasks, they either used large scale private dataset (Qure25k - around 21k samples) for fine-tuning [74] or applied slice-level labels for subject classification [75]. Both cases are not practical for real work clinical application. We want to clarify that our approach, along with recent efforts on building end-to-end 3D foundation models [20, 17, 38], is completely different from either of these two methods to make any direct comparisons. In our case, each scan only has subject-level labels, which constitute a more difficult and clinically realistic task. To be specific, models trained with slice-level labels benefit from substantially more granular supervision—each slice of an imaging volume is explicitly labeled and directly contributes to the training objective. In contrast, our setting only provides a single subject-level label for an entire 3D volume, which is inherently a weaker form of supervision. Our study specifically focuses on the subject-level label setting, which better reflects many real-world clinical scenarios where fine-grained annotations are impractical or unavailable. We therefore emphasize that our contribution lies not in outperforming slice-level supervised methods, but in demonstrating strong performance under the much weaker subject-level supervision paradigm with entire 3D volumetric scan as model input.
Appendix D Additional experiment results
This section provides additional experimental results with more details. Supplementary Figs.˜3 and 4 compares the performance of the foundation model using different numbers of channels and patch sizes, demonstrating that the architecture design of our foundation model is optimal.
Supplementary Fig.˜5 compares our foundation model with a foundation CT model from previous studies under full fine-tuning, Merlin[20], which was pre-trained on abdomen CT scans with corresponding radiology report pairs, CT-FM [38], which was pre-trained on diverse CT scans from different anatomy including head CT. Our model demonstrates superior performance on head CT scans over both models.
Supplementary Fig.˜7 compares our foundation model with Google CT Foundation model [17], CT-FM [38] and Merlin [20] under linear probing. Google CT Foundation were pre-trained with large-scale CT scans from different anatomy. Our model consistently shows improved performance across the board even though our model was pre-trained with less samples than Google CT Foundation.
Supplementary Fig.˜10 presents the volume-to-volume subtype retrieval performance evaluated with Precision@K@K. Our model consistently outperforms CT-FM and Merlin across subtypes. Compared to Google CT, the improvements are more modest, with certain subtypes reaching comparable performance.
Supplementary Figs.˜11 and 12 compares our model with CT-FM under few-shot setting. The result shows our model also presents superior few-shot performance across the board over CT-FM, demonstrating the label efficiency advantage of our model.
Supplementary Fig.˜6 compares the performance on downstream tasks with various supervised tuning methods applied to foundation models pretrained with the MAE and DINO frameworks. Per-pathology comparisons are shown in Supplementary Figs.˜14 and 13. Meanwhile, supplementary Fig.˜15 complements Fig.˜4 in the main article, illustrating the per-pathology performances of foundation models pretrained with different scales of training data.
Supplementary Figs.˜16 and 17 studies the impact of batch effect caused by different CT scan protocols of slice thickness and machine manufacturer. Detailed per-pathology performances are shown in Supplementary Figs.˜18 and 19.
@@ Appendix D Additional experiment results
Supplementary Figure 5: Comparison to previous 3D Foundation Model with Full Finetuning. This plot compares the performance of our model with Merlin [20], CT-FM [38], ResNet3D, our train from scratch and pre-trained model across four datasets. All experiments are done with fine-tuning following the original data pre-processing pipeline of respective models. Our DINO trained model is used in this comparison. Our model demonstrates consistently superior performance across majority of diseases, with the exception of epidural hemorrhage (EDH) in the CQ500 dataset.
@@
Supplementary Figure 7: Performance comparison for linear probing This plot compares our model performance vs. Google CT Foundation model[17], Merlin [20] and CT-FM [38] across all diseases on RSNA and CQ500. Since Google CT Foundation model requires uploading data to Google Cloud (not allowed on our private data) for requesting model embeddings with model weights inaccessible, only public dataset comparison is provided in this study. Similar to other evaluations, we observed that our model outperforms Google CT Foundation model across the board with the only exception on Midline Shift for Google CT Foundation model and EDH for Merlin. Note that CT-FM original data pre-process used width-wise concatenation for different CT windows different from its pre-training input shape, which causes sub-optimal performance when linear probing is performed (not able to adjust backbone parameters for adapting to this specific data input).
@@
Supplementary Figure 8: Performance comparison with 2D and video foundation model This plot shows model performance comparison on our model against MIL using 2D foundation model DINOv3 [40] with attention based multiple-instance learning [39] (simplied as MIL in the plot) and video foundation VJEPA2 [41]. All comparisons are conducted on same image resolution ($\times96$$\times$96) as DINOv3 and VJEPA2 both support dynamic resolutions. Comparing to these two methods, our model shows consistently improvement on AUROC and AP across majority of tasks with significant advantages on inference speed and memory cost (as shown in Supplementary Fig.˜9)
Supplementary Figure 9: Throughput vs. Memory Cost Analysis: in this plot, we perform model efficiency analysis on throughput and memory cost for each of benchmarked models in our study (all models code are taken from official code repository). We use video model terms to standardize input shape naming of each model in this plot where $s$ means height and weight and $f$ means depth (frames for video). The analysis shows that our model with vanilla 3D ViT demonstrates significant advantage in terms of both throughput and memory cost.
(a) RSNA (vs. Google CT)
(b) RSNA (vs. CT-FM)
(c) RSNA (vs. Merlin)
(d) CQ500 (vs. Google CT)
(e) CQ500 (vs. CT-FM)
(f) CQ500 (vs. Merlin)
Supplementary Figure 10: Precision@K@K performance comparison for volume-to-volume retrieval. Solid lines represent our model, dotted lines the compared model, and different colors represent subtypes. Our model maintains competitive performance across all tasks.
NYU Langone
Refer to caption Refer to caption
NYU Long Island
Refer to caption Refer to caption
RSNA
Refer to caption Refer to caption
Supplementary Figure 11: Few-shot performance of different foundation models. We compare different few-shot models performance averaged across diseases for different datasets. The experiment setup follows Fig.˜2 in the main article (evaluated on 8, 16, 32, 64, 128, 256 balanced positive/negative samples). We show our model consistently outperform CT-FM under few-shot setting.
NYU Langone
Refer to caption Refer to caption
NYU Long Island
Refer to caption Refer to caption
RSNA
Refer to caption Refer to caption
Supplementary Figure 12: Few-shot performance of CT-FM. The plots display the per-pathology AUROC and average precision (AP) of the disease detection model under a few-shot learning setting, following evaluation method of Fig.˜2
            NYU Langone
Refer to caption Refer to caption
            NYU Long Island
Refer to caption Refer to caption                RSNA
Refer to caption     Refer to caption                CQ500
Refer to caption     Refer to caption
Supplementary Figure 13: Performance comparison of supervised finetuning methods per pathology on the foundation model trained with DINO. This plot breaks down the average performance across all diseases shown in Supplementary Fig.˜6. The results show that fine-tuning the entire network achieves the best performance in most scenarios. However, linear probing closely approaches finetuning performance for many diseases especially on small or imbalanced dataset, underscoring the capability of our pre-trained models to generate representations that adapt effectively to diverse disease detection tasks.
@@
Supplementary Figure 15: Performance for Different Percentage of Pre-training Samples (Per-Pathology). This plot illustrates label efficiency for individual pathologies using Tukey plots, alongside the average performance across all diseases shown in Figure 1 from main paper. The results indicate that the majority of pathologies show improved downstream performance as the amount of pretraining data increases.
Appendix E Time complexity increase with reduced patch size
@@
Supplementary Figure 16: The 2D projection with t-SNE of CT volume representation extracted from the foundation model. Interestingly, certain subgroups still exhibited slightly better AUCs. For instance, scans with slice thicknesses between 1–4 mm (represented by light green points in the upper panel of Fig.˜16) align with a specialized protocol for CT angiography (CTA), which uses contrast dye to improve diagnosis on particular diseases.